OpenStudy (anonymous):

rationalize the denominator ^3sqrt5y^4/^3sqrt15x^4

OpenStudy (turingtest):

someday I hope you will learn to use the equation feature...\[\sqrt[3]5y^4\over \sqrt[3]15x^4\]or\[\sqrt[3]{5y^4}\over \sqrt[3]{15x^4}\]???

OpenStudy (anonymous):

they both look the same to me

OpenStudy (turingtest):

are the variable under the radical or just the numbers?

OpenStudy (anonymous):

its written how you have it written with a thin line not bold

OpenStudy (turingtest):

I don't understand, is it the first way I posted it or the second?

OpenStudy (anonymous):

yes the 1rst

OpenStudy (turingtest):

\[{\sqrt[3]5y^4\over \sqrt[3]{15}x^4}={y^4\over \sqrt[3]3x^4}*{\sqrt[3]{3^2}\over \sqrt[3]{3^2}}={\sqrt[3]9y^4\over3x^4}\]sorry that took a while, I've never actually rationalized a non-square root before and had to figure it out.

OpenStudy (anonymous):

its okay thank you :D